# PyTomography: A Python Library for Medical Image Reconstruction Luke Polson1,2, Roberto Fedrigo1,2, Chenguang Li1,2, Maziar Sabouri1,2, Obed Dzikunu2,3, Shadab Ahamed1,2, Nicolas Karakatsanis4, Arman Rahmim1,2,3,5,∗, Carlos Uribe2,5,6,∗

### v3

@@ II.B. Software Architecture
    Data. The standard data type used to represent images and raw projection data is the tensor class of PyTorch.
@@
    Metadata. Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
@@
    Transforms. Transform classes are used as utilities to model various features of the imaging process, such as attenuation and image blurring in SPECT. They are commonly used as components in a system matrix. Specifically, transforms are defined by the mathematical property that they are linear endomorphisms.
@@
    Projectors. Unlike transforms, projectors are mathematical operators that map between different vector spaces. The system matrix, which maps from object space to projection space and represents a model of the imaging system, is an example of a projector. While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries.
  5. 5.
    Likelihoods. Likelihoods specify the likelihood function that characterizes the acquired data. Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$.
  6. 6.
@@
    Algorithms. Algorithms are used for image reconstruction, and are designed to interface with all imaging modalities. Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks.
A visual representation of the software library along with sample code representing a SPECT reconstruction use case is shown in Figure 1. Metadata is first obtained via the raw vendor data stored in the path represented by the variable “file_NM”. The metadata is used as input to a “SPECTSystemMatrix” class, which also takes as input attenuation, PSF, and field of view cutoff transforms (creation of transforms not shown). The generated system matrix is used to build a likelihood instance, which also takes in as input the photopeak (creation not shown) and scatter as an additive term (creation not shown). The likelihood is then used to initialize an OSEM class for reconstruction; a reconstructed object is subsequently obtained using 4 iterations and 8 subset.
Figure 1: Basic overview of the workflow of PyTomography (top) with code representing a specific SPECT reconstruction use case (bottom). Sample classes and implementations are shown; these do not comprise the full set of classes available in the software.
II.C. Examples
@@
Multiple examples are explored to highlight the capabilities of PyTomography. At present, reconstruction of parallel collimator SPECT and LM PET data is demonstrated, but it should be emphasized that these are purely intended as example use cases for particular imaging modalities. All computations were performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a TeslaV100 GPU. Python scripts and notebooks used to obtain these results can be found at https://github.com/qurit/PyTomography_paper_code.
II.C.1. Monte Carlo SPECT
Two use cases are explored for reconstruction of Monte Carlo generated SPECT image data from the SIMIND Monte Carlo simulation program26. The system matrices used in each case correspond to equation 1, with each employing attenuation modeling and PSF modeling. The triple energy window technique27 is used for scatter correction.
[…]
Table 4: Reconstruction parameters used for SPECT reconstruction in this paper. Neighbourhood size corresponds to the kernel size in KEM, and prior neighbourhood for BSREM. Nearest neighbours were obtained from attenuation maps.
Algorithm OSEM BSREM KEM
Subsets 8 8 8
Neighbourhood Size - $3\times 3\times 3$ $5\times 5\times 5$
$k$ Nearest Neighbours - 8 40
Kernel - - Eq. 5 of Vuohijoki et. al. 6
Prior - RDP-AP -
[…]
II.C.2. Clinical SPECT
This example highlights SPECT reconstruction on publicly available data from the Deep Blue data repository31 that was used for the SNMMI 177Lu dosimetry challenge32, 33. Patients received radiopharmaceutical therapy with 177Lu-DOTATATE for neuroendocrine tumours (4 cycles of 7.4 GBq/cycle administered every 8 weeks); scans were taken at 4 different time points in the week following a therapeutic dose. Images consisted of 120 projections of shape $256\times 256$ with resolution $2.4~{}\text{mm}\times 2.4~{}\text{mm}$. The system matrix employed attenuation, PSF, and finite FOV modeling. Reconstructions using OSEM, BSREM, and KEM were performed for 50 iterations each; the TEW technique was used to correct for scatter in all cases.
II.C.3. Monte Carlo PET
This example demonstrates (i) the capabilities of LM TOF PET reconstruction on data generated by the GATE Monte Carlo simulation program and (ii) the DIP Recon technique. The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps. For simplicity, data acquisition in the simulation ignored dead-time effects and only considered true primary coincidence events; coincidences due to phantom-scatter and randoms were ignored. The digital phantom used in simulation corresponded to an ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34; the full simulated dataset was representative of a 3 min scan and consisted of $556$ million coincidence events. Data was reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. Reconstruction was performed with normalization correction in 3 different ways:
@@ II.C.3. Monte Carlo PET
    The full dataset (556 million events) was reconstructed using OSEM (4it, 20ss) using the EM algorithm
@@
    6.3% of the data set (35 million events) was reconstructed using OSEM (80it, 1ss) using the EM algorithm, representative of a low-count scenario.
@@
    6.3% of the dataset (35 million events) was reconstructed using the DIPRecon algorithm. The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the previous low count reconstruction. The neural network was a UNet35 with similar architecture to the original DIP paper 8; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
III. Results
The time required for reconstruction in the first SPECT use case was 412.8 s (3.44 s/iteration) for OSEM, 434.8 s (3.62 s/iteration) for BSREM, and 601.8 s (5.02 s/iteration) for KEM; sample reconstructions with corresponding RC curves are shown for each algorithm in Figure 2. BSREM clearly produced the least noisy images, and had the best RCs for the lungs, liver, kidneys, and bladder. KEM produced an image estimate with an intermediate level of noise, and had the best recovery coefficient for the salivary glands. OSEM yielded the noisiest images, and had the worst recovery coefficients in all regions except the salivary glands. Despite the high performance of BSREM, it was unable to accurately predict the regions of low uptake within the liver visible in the ground truth.
Figure 2: Reconstructions of full body 177Lu-PSMA-617 Monte Carlo projection data; ground truth phantoms were obtained using the XCAT software and SPECT simulation was achieved using the SIMIND Monte Carlo program. Left: Central coronal slices corresponding to truth, OSEM (120it/8ss), BSREM (120it/8ss), and KEM (120it/8ss). Right: Recovery coefficient vs. iteration number for sample organs; line widths are representative of error estimates based on reconstructions of 5 independent noise realizations.
Sample reconstructions and RC curves for the SPECT motion correction use case are shown in Figure 3. The lesions appear sharper when motion correction is used; motion correction applied during reconstruction via Equation 3 produced lesions of higher uptake than when correction was applied post-reconstruction. This is quantitatively observed in the RC curves of Figure 3, where motion correction applied during reconstruction tends to produce the highest RCs throughout all iterations. After 100 iterations, the RCs were larger when MC was used during reconstruction than when it was used post reconstruction ($p<10^{-6}$).
Figure 3: Reconstructions of full body 177Lu-PSMA-617 Monte Carlo projection data with breathing motion simulated; ground truth phantoms were obtained using the XCAT software and SPECT simulation was achieved using the SIMIND Monte Carlo program. Reconstruction was performed with OSEM (8ss) up to 100 iterations using attenuation/PSF modeling and TEW scatter correction; images shown correspond to reconstruction after 20 iterations. From left to right: ground truth phantom, reconstructed image without MC, reconstructed image with MC applied after reconstruction, reconstructed image with MC applied during reconstruction, and lesion recovery coefficients as a function of iteration number. Green boxes are drawn around the lesion in the liver.
The time required to reconstruct the clinical SPECT patient data was 464.7s (9.29 s/iteration) for OSEM, 453.1s (9.06 s/iteration) for BSREM, and 1825s (36.50 s/iteration) for KEM. Central coronal slices are shown in Figure 4; the patient had a region of high uptake value in the liver, indicative of a liver lesion. BSREM and KEM appear to confine activity within the boundary of the kidneys, with BSREM exhibiting slightly more of this behaviour. In addition, OSEM produced the most observed noise in the liver. A common artifact of both BSREM and KEM, however, is the noise in the liver lesion; use of anatomical nearest neighbours may cause the lesion activity noise profile to resemble the CT noise profile. The mean predicted counts in the kidneys were 1.98 (OSEM), 2.07 (BSREM), and 2.02 (KEM); these relative proportions are consistent with the RCs of Figure 2, suggesting that BSREM may yield the most accurate recovery coefficients for the kidneys.
Figure 4: Coronal slices of activity concentration (colored) and Hounsfield Unit values (greyscale) reconstructed using OSEM (50it/8ss), BSREM (50it/8ss), and KEM (50it/8ss). Data corresponds to the first time point of patient 4 from the “Lu-177 DOTATATE Anonymized Patient Datasets: Multi-Time Point Lu-177 SPECT/CT Scans” dataset31.
Reconstructing the LM TOF PET GATE output data required multiple steps. While initial processing of the ROOT listmode data is time consuming, PyTomography can save this data in an accessible format for reuse.
  • Opening the ROOT list mode data ($\sim$ 50 GB) of the normalization scan and computation of the normalization factor $\eta$ (approximately 50 minutes). When the normalization factor is saved, it requires approximately 5 s to read.
  • Opening the ROOT list mode data ($\sim$ 74 GB) of the phantom scan (approximately 1 hour). When saved in an accessible format, reading requires approximately 5 s.
  • Initial configuration of the system matrix ($7.8$ s).
  • Reconstruction of the PET listmode data. The high count scenario using OSEM (4it 20ss) required 185.6 s (46.4 s per iteration), the low count scenario using OSEM (80it 1ss) required 201.0 (2.5 s per iteration), and the low count scenario using DIPRecon required 3135.6 s (31.3 s per iteration)
Corresponding phantom reconstructions are shown in Figure 5; bias and noise in grey and white matter is shown as a function of iteration number in Figure 6. Due to finite resolution of the imaging system causing activity “leakage”, activity quantification in the grey matter region was consistently underestimated while activity quantification in the white matter region was consistently overestimated. The DIPRecon on the low count data had similar noise levels in grey matter to OSEM on the high count data, demonstrated both qualitatively and quantitatively in each Figures 5 and 6 respectively.
Figure 5: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (20it/4ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction.
@@ III. Results
Figure 6: Variation in absolute bias and standard deviation throughout iterative reconstruction corresponding to reconstructed images of Figure 5. Subplots correspond to the grey matter region (left) and white matter region (right). For the low count rate scenario, points are plotted every 10 iterations. For the low count rate scenario with OSEM, the first point corresponds to the 10th iteration. The bias was negative in grey matter and positive in white matter.
IV. Discussion
@@
This technical report presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data. The short times required for reconstruction using this tool can enable researchers to perform extensive studies on phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms. Validation of PyTomography against alternative reconstruction libraries is shown in the appendix.
Example applications in 177Lu-SPECT imaging, which are relevant for currently FDA approved radiopharmaceutical therapies, explored the use of the OSEM, BSREM, and KEM, reconstruction algorithms. It should be emphasized that the purpose of this technical report was not to rank the different algorithms, but rather to demonstrate the capabilities of the software library. While some evidence was shown to favour BSREM over OSEM and KEM for kidney and liver dosimetry based on better recovery coefficients (Figure 2), a truly comprehensive study that aims to quantitatively compare algorithms should involve (i) multiple XCAT activity and anatomical configurations, (ii) inclusion of patient motion, (iii) effect of misalignment between SPECT and CT images, (iv) different hyperparameters in each reconstruction algorithm and (iv) should seek to use real patient data to validate any phantom studies. In addition, recovery coefficients alone are not sufficient to rank algorithms. Studies should also consider source-to-background ratios as well as error analysis based on intra- and inter-patient variabilities, and clinical task performance. While the presented software framework enables such studies to be performed, it is out of scope for this publication.
While the RDP-AP prior was featured heavily in this study, the framework of PyTomography is designed to foster the development and validation of other novel prior functions as well. In addition, while KEM was featured using anatomically-based basis functions, the library permits use of any external image for construction of the basis functions. Different priors and basis functions may also have significant implications for (i) quantitative PET/SPECT imaging in dosimetry, and (ii) qualitative observer based studies in lesion detection and classification.
The motion correction example considered a simplified case of 4 discrete phases where motion correction transforms were obtained using ground truth attenuation maps. In practice, however, breathing motion is continuous, and motion correction transforms will need to be derived using alternative source of data. The scope of this technical report is merely to demonstrate that PyTomography contains the functionality for incorporating motion correction.
@@
Example applications in 18F-FDG PET imaging demonstrated the capabilities of LM reconstruction using the OSEM and DIPRecon algorithms. As demonstrated qualitatively in Figure 5 and quantitatively in Figure 6, reconstruction of the low count data using DIPRecon was able to reduce noise in the grey matter regions while preserving cortical structures. Once again, the purpose of this paper is merely to demonstrate the capabilities of PET LM reconstruction and use of the DIPRecon algorithm provided in the software library; extensive research on the validity of the DIPRecon method should seek to explore, for example, the effects of different neural network architectures, hyperparameters, and misalignment between the PET/MRI images.
@@
While this paper has demonstrated implementations for parallel collimator SPECT and LM TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, sinogram-based PET reconstruction, and various forms of computed tomography (CT) and magnetic resonance (MR) imaging. As displayed in Figure 1, projectors for different imaging modalities can be implemented using a variety of python based tools: it is hoped that the structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems. Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own data. The immediate goals for future development in PyTomography at the time of publication include:
@@
    Incorporation of scatter and random estimation in PET reconstruction.
@@
    Development of a 3D Slicer36 extension.
In summary, we hope that we have made a compelling case for the benefit that this tool can bring to the medical imaging community. It can be used as a framework for studies that require fast reconstruction times, such as those that explore a wide range of reconstruction hyperparameters. It is well validated and may be used for accurate and reproducible quantitative imaging in both PET and SPECT; we particularly envision applications in the field of theranostics. A variety of developments will assist PyTomography in reaching the full potential of precision medicine. We thus highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.
V. Conclusion
This work describes the python library PyTomography and highlights specific use cases in SPECT and PET imaging. The software architecture facilitates the development of different imaging modalities that can all interface with the same reconstruction algorithms. The purpose of PyTomography is to create a well-documented and computationally-efficient library for medical image reconstruction, where novel system modeling and reconstruction techniques are implemented, shared, and explored by experts in the community.
VI. Appendix
VI.A. Relevant Links
Links to the various webpages of PyTomography are contained in Table 5: GitHub contains the source code, Readthedocs contains all relevant documentation, PyPI is the host of the built source files, and Discourse is a community forum for users and contributors.
Table 5: Relevant Webpages
Website Link
Github https://github.com/qurit/PyTomography
Readthedocs https://pytomography.readthedocs.io/en/latest/
PyPI https://pypi.org/project/pytomography/
Discourse https://pytomography.discourse.group/
VI.B. Validation of PyTomography
This section contains two examples used to validate PyTomography against other reconstruction software:

### v4

@@ 2.2 Software architecture
    Data/Metadata. The standard data type used to represent images and raw projection data is the tensor class of PyTorch; data can be stored on CPU or GPU. Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
@@
    Projectors. Projectors are mathematical operators that map between object and projection space. The system matrix is an example of a projector that models a particular imaging system. While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries. Projector subclasses inherit from a generic “SystemMatrix“ class and implement a “forward”, “backward”, and “compute_normalization_factor” method. Projectors used in subset based algorithms should additionally implement “set_n_subsets”, “get_projection_subset” and ”get_weighting_subset”; the second method takes in projection data and returns subset data based on how the projector defines subsets, and the third gets the corresponding weight ($c^{n}$) for scaling preconditioners.
@@
    Likelihoods. Likelihoods specify the likelihood function that characterizes the acquired data. Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$. Likelihoods inherit from a generic “Likelihood” class and should implement the “compute_gradient” method. Higher order derivatives of the likelihood may also be defined as seperate methods within the class.
@@
    Algorithms. Algorithms are used for image reconstruction, and are designed to interface with all imaging modalities. Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks. Unlike projectors and likelihoods, algorithms do not inherit from a generic base class. In particular, however, linear PGAAs inherit from a generic “LinearPreconditionedGradientAscentAlgorithm” class and simply need to implement the “_linear_preconditioner_factor” and optionally reimplement the “__init__” method to take in required arguments.
PyTomography also has an input/output (I/O) module for reading various file formats; it currently supports SIMIND, DICOM (SPECT/CT), GATE (PET), and HDF5 listmode (PET) formats. Data reading consists of loading the generated files and constructing the appropriate data/metadata instances within the PyTomography framework. The I/O module is also used to read internal files of the library, such as coefficients to convert Hounsfield Units to attenuation coefficients, files containing SPECT collimator information, and files containing PET scanner geometry.
Novel development of a projector with associated data sources should thus consist of the following (i) implementation of metadata classes, (ii) I/O code to load data/metadata from the data sources, and (iii) development of a projector to interface with the data/metadata. Metadata class instances may be initialized by various data sources; for example, the same “SPECTProjMeta” class is initialized from SIMIND or DICOM I/O data loading functions, even though they represent different data sources. Metadata is thus linked to a modality and not to a specific data source.
Deployment of the software was designed to maximize accessibility. In addition to placing the backend code on GitHub, the library was also uploaded to the Python Package Index (PyPi) so users can easily install the library using the simple “pip install pytomography” terminal command. A readthedocs documentation website was created that also contains a comprehensive set of tutorials in jupyter notebook format. While many of the tutorials are aimed at training end users, there also exist developer tutorials for implementing new projectors in the software. A discourse website was developed for two purposes: (i) to allow end users to ask questions about how to use the software and (ii) to permanently store and make public the questions/answers as they may be useful for other users of the library.
Figure 1: Basic overview of the workflow of PyTomography when using statistical algorithms (left) with code representing a specific SPECT reconstruction use case (right). Shown are the minimum required inputs to the “__init__” function. the minimum required methods for each class, and sample subclasses available in the library. The “…” represents additional arguments that may be present in subclasses. In the SPECT use case, for example, the additional input argument “obj2obj_transforms” is implemented to take in operations that implement attenuation and PSF modeling. Not shown in the sample code: importing required libraries, loading the photopeak/scatter data, creating the attenuation/PSF transforms, and specifying the filepath “file_NM”.
3 Illustrative examples
Multiple examples are explored to highlight the capabilities of PyTomography. Specifically, reconstruction of parallel-collimator SPECT and LM-TOF PET data is demonstrated, but it should be emphasized that these are only intended as example use cases for particular imaging modalities. Validation against other software libraries is also shown in the Appendix, which contains (i) SIMIND simulated SPECT data reconstruction validated against STIR, (ii) DICOM SPECT data (obtained on a GE Tandem Discovery 670 Pro) validated against MIM’s SPECT reconstruction functionality, and (iii) GATE PET data reconstruction validated against CASToR. We chose to validate DICOM reconstruction against MIM, and not the Siemens internal reconstruction software, because it represents an external library with access to the same information as PyTomography.
@@ 3 Illustrative examples
All computations were performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a TeslaV100 GPU. Python scripts and notebooks used to obtain these results can be found at https://github.com/qurit/PyTomography_paper_code.
3.1 SPECT
This section highlights SPECT reconstructions using available reconstruction algorithms in PyTomography, demonstrating differences in quantitative metrics between different algorithms. Two examples are considered: (i) reconstruction of Monte Carlo SPECT data generated via SIMIND, and (ii) reconstruction of clinical SPECT data. In the clinical case, the patient received radiopharmaceutical therapy with 177Lu-PSMA for prostate cancer and was imaged 3 hours post injection; the XCAT phantom used as input for SIMIND was configured such that the kidney, liver, and lesion activity concentrations roughly approximated the clinical patient. In each case, the data correspond to acquisition on a GE Tandem Discovery 670 Pro SPECT system (MEGP collimators) with 60 projections (25 s per projection). The clinical data consists of acquisition at two bed positions with projections of shape $128\times 128$, while the simulated data consists of acquisition at one bed position with projections of shape $128\times 240$. The two lesions inserted in the XCAT phantom were (i) based on an observed bone metastasis in the sternum of the clinical patient, and (ii) placed in the liver. In each case, the system matrix employed attenuation and PSF modeling, and the likelihood function considered an additive scatter term obtained via the triple energy window technique.
The reconstruction algorithms considered were OSEM (4it/6ss), BSREM-RDP (40it/6ss), BSREM-RDPAP (40it/6ss), and KEM (10it/6ss) [’it’ for iterations; and ’ss’ for subsets]; algorithms using RDP used the hyperparameters $\beta=0.3$, $\gamma=2$, RDPAP only considered the 8 nearest neighbours based on the attenuation map, and KEM used Eq. 5 of Vuohijoki [6] for its filter. The time required for reconstruction of one bed position of the clinical data was 1.07 s/it (OSEM), 0.83 s/it (BSREM-RDP), 0.86 s/it (BSREM-RDPAP), and 1.07 s/it (KEM). The time required for reconstruction of the simulated data was 1.32 s/it (OSEM), 1.13 s/it (BSREM-RDP), 1.21 s/it (BSREM-RDPAP), and 1.52 s/it (KEM).
For the clinical data, the left kidney was segmented using the CT-based TotalSegmentator model [24], and the bone metastasis in the sternum was segmented by first segmenting the lesion from the OSEM reconstruction, and then applying a threshhold mask greater than 0.150.150.15~{}0.15cm-1 so that only voxels within the bone were considered. A spherical mask was used in the liver for both the simulated and clinical data for measurement of noise in each reconstruction algorithm; noise was quantified as a percent by computing the standard deviation of uptake divided by the mean of uptake in that region. Reconstructions of the clinical and simulated data are shown in Figures 2 and 3 respectively.
Figure 2: Top: Maximum intensity projections of reconstructed clinical SPECT data; the leftmost subplot corresponds to reconstruction using the MIM reconstruction software library. Three annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL. Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
Figure 3: Top: Maximum intensity projections of reconstructed simulated SPECT data; the leftmost subplot corresponds to the ground truth XCAT phantom used for simulation in SIMIND. Four annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL, and (iv) the location of the liver metastasis (purple arrow) with the corresponding activity concentration in kBq/mL. Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
3.2 PET
PyTomography is capable of reconstructing real and simulated PET data in either sinogram mode or listmode. This example uses GATE Monte Carlo data and demonstrates the following capabilities of the library in PET reconstruction:
@@ 3.2 PET
    LM-TOF reconstruction
@@
    Random estimation via delayed coincidences and scatter estimation using the TOF single scatter simulation (TOF-SSS) technique
@@
    Use of the DIPRecon algorithm to enhance reconstruction
The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities. The digital phantom used for simulation in GATE consisted of an ultra-high resolution 18F-FDG PET image, a T1-weighted MRI brain image, and a corresponding attenuation map [25]. The total uptake of 15.66 MBq was representative of a typical 18F-FDG distribution 1 hour post injection. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. In all cases, simulated data were reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm; 5 TOF bins were used with a total range of 30 cm. Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction. The “scatters” contribution for each listmode event was estimated using the TOF single scatter simulation (SSS) capabilities of the library, based on Watson [26]. The PET image used as a proxy for SSS was a reconstruction of the data via LM-TOF OSEM (40it, 1ss) without scatter estimation. Simulation and reconstruction were then performed in the following three ways:
  1. 1.
    A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan. The data was then reconstructed using LM-TOF OSEM (80it, 1ss) with attenuation, random, and scatter modeling. PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 4 mm.
  2. 2.
    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed.
  3. 3.
@@
    A low count reconstruction was performed without PSF modeling and used in the DIPRecon algorithm; no PSF modeling was used because (i) the DIP performs better on images with high frequency noise, and (ii) the DIPRecon algorithm already has noise reduction properties. The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the low count reconstruction. The neural network was a UNet [27] with similar architecture to the original DIP paper [8]; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
The time required for the complete TOF-SSS estimation was 665 s (high count) and 577.5 s (low count), the time required for the OSEM reconstructions were 3.31 s/it (high count) and 1.18 s/it (low count), and the time required for the deep image prior reconstruction was 289.1 s (initial fit) and 21.53 s/it (reconstruction). Corresponding phantom reconstructions, as well as bias and noise in grey and white matter as a function of iteration number, are shown in Figure 4.
@@
Figure 4: Left: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25]. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (80it/1ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction. Right: Variation in absolute bias and standard deviation throughout iterative reconstruction. Subplots correspond to the grey matter region (upper) and white matter region (lower). For the low count rate scenario, points are plotted every 10 iterations. For the low count rate scenario with OSEM, the first point corresponds to the 10th iteration. The bias was initially positive and became negative for all three cases in white matter.
4 Impact
@@
In this work, we have motivated and presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data. The short times required for reconstruction using this tool can enable researchers to perform extensive studies on digital/physical phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms. To demonstrate (i) the impact of the software and (ii) the extent to which new research questions can be pursued, we will discuss each of the illustrative examples.
A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library. For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP. Furthermore, the simulated data prediction from BSREM-RDPAP (807.02.17807.02.17807.02.17~{}807.02.17kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration. For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise. BSREM-RDPAP has a lower noise (6.9%) than OSEM (13.3%) but with higher uptake in the kidney (807.02 kBq/mL for BSREM-RDPAP vs. 755.07 kBq/mL for OSEM) for the simulated data; BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction. For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown on Figures 2 and 3. These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (217.13217.13217.13~{}217.13kBq/mL for BSREM-RDPAP and 216.15216.15216.15~{}216.15kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (198.88198.88198.88~{}198.88kBq/mL). For the liver lesion in the simulated data, however, use of the anatomical reconstruction algorithms decreased the activity concentration predictions (194.04194.04194.04~{}194.04kBq/mL for BSREM-RDPAP and 231.31231.31231.31~{}231.31kBq/mL for KEM) compared to OSEM (241.08241.08241.08~{}241.08kBq/mL). This occured because there is less anatomical difference in the surrounding tissue in the liver metastasis compared to the bone metastasis, so the smoothness prior merely blurred and reduced the activity concentration. These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry. While this example focused primarily on organ and lesion dosimetry, PyTomography users will be to assess the merits of various reconstruction algorithms on any task specific metric.
@@
Example applications in 18F-FDG PET imaging demonstrated the capabilities of LM TOF SSS modeling and LM-TOF reconstruction using the OSEM and DIPRecon algorithms. As apparent in Figure 4, reconstruction of the low count data using DIPRecon was able to reduce noise in the grey matter regions while preserving cortical structures. While use of the DIPRecon on the low count data resulted in a slight increase in bias in the grey matter region, the corresponding noise bias curve demonstrates less noise than the high count reconstruction using OSEM. Extensive research on the validity of the DIPRecon method may seek to explore, for example, the effects of different neural network architectures, hyperparameters, and misalignment between the PET/MRI images. Since PyTomography directly interfaces with PyTorch, the process of designing and testing neural networks and optimization procedures for the DIPRecon method is straightforward. Users who wish to test different network/optimization configurations are encouraged to view the DIPRecon tutorial available on the documentation website.
@@
While this paper has demonstrated implementations for parallel collimator SPECT and LM-TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, various forms of computed tomography, (CT), magnetic resonance (MR) imaging, and compton camera (CC) imaging. As displayed in Figure 1, projectors for different imaging modalities can be implemented using a variety of python based tools: it is hoped that the structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems. Implemented code does not necessarily need to be written directly in python; the parallelproj library used for PET modeling, for example, uses a CUDA C backend for its calculations. This flexibility permits contributions from users familiar with a variety of coding languages. Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own data. The immediate goals for future development in PyTomography at the time of publication include:
@@
    Expanded support for clinical PET data.
@@
    Development of a 3D Slicer [28] extension.
In summary, we hope that we have made a compelling case for the benefit that this tool can bring to the medical imaging community. It can be used as a framework for studies that require fast reconstruction times, such as those that explore a wide range of reconstruction hyperparameters. It may also be used by developers seeking to implement new system matrices and likelihoods for their own imaging systems. We highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.
5 Conclusions
This work describes the python library PyTomography and highlights specific use cases in SPECT and PET imaging. The software architecture facilitates the development of different imaging modalities and likelihoods that can all interface with the same reconstruction algorithms. The goal of this research was to create a simple-to-use and computationally-efficient library for medical image reconstruction, where system modeling and reconstruction techniques are implemented, shared, and explored by experts in the community.
Acknowledgements
Appendix A Validation of PyTomography
This section contains three examples used to validate PyTomography against other reconstruction software: